OpenStudy (anonymous):

**WILL MEDAL** A cylindrical cardboard mailing tube has a diameter of 90 mm. and a height of 600 mm. A total of 20% of the tube’s capacity is filled with spherical packing foam. Each foam sphere has a radius of 15 mm. Approximately how many pieces of foam are in the tube? Use 3.14 to approximate pi and express your answer as a whole number.

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@aaronq

OpenStudy (amistre64):

find 20% of the volume of the cylindar and divide it by the volume of a sphere is my thought

OpenStudy (amistre64):

whats our formulas for these things?

OpenStudy (anonymous):

My mom and I tried to find the answer and we came out to 210 or 659

OpenStudy (amistre64):

the answer will become apparent at the end of the process; lets make sure we are using the correct stuff first what are the formulas we will need to use? volume of cyl, and volume of sphere ...

OpenStudy (anonymous):

Ok volume of a sphere is 4/3*pi*r^3

OpenStudy (amistre64):

good so far, and cylindar?

OpenStudy (anonymous):

the volume of a cylinder = pi* r^2* h r is the radius and h is the height

OpenStudy (amistre64):

good, and we are going to have 2 different radius measures .. what are they?

OpenStudy (anonymous):

3in and 4.5in?

OpenStudy (amistre64):

20% of the cylindar divided by the volume of a sphere .20 cylindar ---------- sphere .20 pi R^2 ------------ 4/3 pi r^3 pi/pi = 1 so the factors cancel .20 R^2 ---------- 4/3 r^3 .20/(4/3) = .20(3/4) = .05(3) = .15 so this simplifies to this as a formula for us: .15 R^2/r^3

OpenStudy (amistre64):

daims are 90 and 15 ... what are the radiuses?

OpenStudy (anonymous):

45 and 7.5

OpenStudy (amistre64):

good, i noticed i dropped the h someplace along the way ... h is 600 .15 h R^2/r^3 thats better so we just fill in .15 (600) (45)^2/(7.5)^3

OpenStudy (anonymous):

96.426

OpenStudy (amistre64):

hmm, you might need new batteries in your calculator :) http://www.wolframalpha.com/input/?i=.15+%28600%29+%2845%29%5E2%2F%287.5%29%5E3

OpenStudy (anonymous):

Um i dont think so lol I used an online one

OpenStudy (anonymous):

So 432?

OpenStudy (amistre64):

then your inputs were mistyped. either way ... we are getting an upper limit of 432. yes

OpenStudy (anonymous):

So do we multiply 432 and pi?

OpenStudy (amistre64):

no, pi was a common factor that canceled out. there is no need to work in pi

OpenStudy (anonymous):

Oh ok so the answer is 432

OpenStudy (amistre64):

yes, but how good of an approximation that is i wouldnt know how to tell we know that at most, 432 spheres is equal to the volume required .. but if you stuff them in there their shape leaves gaps ... so they fill MORE than 20%

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

The answert was 54

OpenStudy (amistre64):

how is the in any way a possibility

OpenStudy (amistre64):

if we assume a cube instead of a sphere? .20(3.14)(45)^2(600)/15^3 thats 226

OpenStudy (amistre64):

did you mis type the information any?